OpenStudy (anonymous):

Use a graphing calculator to determine which equation for the line of regression, Pearson product-moment correlation value (r), and predicted value match the data in the chart. The chart below shows the number of hours a student spent studying for a math test and the score the student earned on the test.

OpenStudy (anonymous):

I'm posting the work

OpenStudy (anonymous):

Hours Spent Studying | 0.5|1.25| 1.5 |1.75 |2 | 2.5 |3 |3.5 |4 | Test Score,,,................ | 90 | 75| 60 | 83 |75| 80| 99| 97| ? | A) y = 6.4x + 69.5, r = 0.48 Predicted test score for 4 hours of study equals 95. B) y = 6.1x + 74.5, r = 0.48 Predicted test score for 4 hours of study equals 99. C) y = 5.1x + 74.5, r = 0.88 Predicted test score for 4 hours of study equals 95. D) y = 7.4x + 69.5, r = 0.88 Predicted test score for 4 hours of study equals 99.

OpenStudy (anonymous):

Do you know how to turn the (r) on on your calculator?

OpenStudy (anonymous):

I don't have a calculator yet.... I need a ti-83 or 84 and my mom hasn't bought one yet :(

OpenStudy (anonymous):

The online graphing calculators are hard to use too...

OpenStudy (anonymous):

Then you can't do the problem. Google "virtual TI 84" and you should be able to download one

OpenStudy (anonymous):

I'm so stuck please help someone!

OpenStudy (helder_edwin):

u can use excel or any ohter spreadsheet program.

OpenStudy (anonymous):

but how do I plot it in excel?

OpenStudy (anonymous):

And thanks so much btw

OpenStudy (helder_edwin):

give me a couple of min.

OpenStudy (anonymous):

okay thanks so much you are the best!

OpenStudy (helder_edwin):

here is the excel file

OpenStudy (anonymous):

Thanks! SOOO Much!

OpenStudy (helder_edwin):

u r welcome